### v4

# Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling

Abstract
Large language models (LLMs) are prone to hallucination stemming from misaligned self-awareness, particularly when processing queries exceeding their knowledge boundaries. While existing mitigation strategies employ uncertainty estimation or query rejection mechanisms, they suffer from computational efficiency and sacrificed helpfulness. To address these issues, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework, integrating fast and slow reasoning systems to harmonize reliability and usability. The framework first employs a fast-thinking model to generate confidence-labeled responses, enabling immediate utilization of high-confidence outputs, whereas uncertain predictions trigger a slow refinement model for accuracy improvement. To align model behavior with our proposed object, we propose a hybrid training pipeline, enhancing self-awareness without degrading task performance. Evaluations on dialogue state tracking tasks demonstrate that EKBM achieves superior model reliability over uncertainty-based baselines. Further analysis reveals that refinement substantially boosts accuracy while maintaining low computational overhead. The framework establishes a scalable paradigm for deploying reliable LLMs in error-sensitive applications, effectively balancing accuracy and practical utility.
1 Introduction
Recently, large language models (LLMs) have demonstrated impressive text generation capabilities (Abdullah et al., 2022). However, LLMs are susceptible to hallucinations, where generated content misaligns with context or factual information (Zhang et al., 2023). Hallucinations can be particularly detrimental in applications with low tolerance for error, eroding user trust in LLM reliability.
Extensive works have been devoted to mitigating hallucinations in LLMs (Tonmoy et al., 2024). Hallucinations often arise from misalignments between an LLM’s outputs and its intrinsic knowledge boundaries, particularly when when addressing queries beyond its expertise (Li et al., 2024b). Improving the model’s self-awareness—its capacity to accurately assess its own knowledge boundaries and outputs correctness, can effectively mitigate this misalignment. Uncertainty-based methods estimate model confidence and indirectly reflect the model’s self-awareness (Huang et al., 2024), facing challenges of inherence to the model’s endogenous capabilities, frequently sufferring from high computational costs or instability across thresholds. Alternative strategies emphasize aligning knowledge boundaries with actions by rejection of out-of-scope queries (Xu et al., 2024b), avoiding mistakes but sacrificing helpfulness—a trade-off that undermines model utility particularly in complex, multi-step tasks where partial correctness remains valuable.
To address these limitations, we propose a novel alignment objective that enhances model awareness of its knowledge boundaries, enabling models to respond more effectively by explicitly distinguishing between high-confidence (“sure”) and low-confidence (“unsure”) outputs. A reliable system should deliver near-perfect accuracy for “sure” predictions and provide helpful information in the “unsure” category, serving as a form of soft rejection, appropriately managing user expectations regarding accuracy. Building on this concept, we propose the Explicit Knowledge Boundary Modeling (EKBM) framework. EKBM integrates both Fast and Slow systems: a fast-thinking model generates responses annotated with confidence labels, allowing for immediate use of high-confidence outputs, while a slow-thinking refinement model engages in deliberate reasoning to improve low-confidence predictions. Crucially, EKBM does not merely filter uncertain outputs but
leverages them as opportunities for improvement, ensuring both precision and recall. The collaboration between these two systems strikes a balance between reliability and efficiency, with the fast-thinking model’s self-awareness being crucial for the system’s effective functioning.
To operationalize this paradigm, we design a training pipeline that aligns the model’s capability boundaries and improves self-awareness. Evaluations on dialogue state tracking (DST) tasks demonstrate that our method effectively and reliably distinguishes the confidence levels of model output compared to baseline methods. Additionally, we develop an automatic refinement model for “unsure” outputs. Experiments show that the EKBM system achieves superior overall performance compared to traditional approaches that directly tune models for the task.
Our contributions are summarized as follows:
We present the EKBM framework, integrating fast-thinking assessment with slow-thinking refinement, along with metrics for evaluating self-awareness and reliability.
We develop a comprehensive training pipeline that enhances LLM self-awareness, yielding more reliable models.
We perform extensive experiments to demonstrate the effectiveness and scalability of the EKBM framework, providing valuable insights for future research.
2 Problem Formulation
2.1 Alignment for Reliability
To enhance the reliability of LLMs, a prevailing method involves aligning their performance with high-quality training data. This alignment aims to achieve the following objective:
where a scoring function ensures that the score of a correct response to an input prompt is higher than that of an incorrect response .
Recent studies have introduced rejection as a mechanism to help the model differentiate between answerable (in-boundary) and unanswerable (out-of-boundary) queries, where model is encouraged to directly reject the unanswerable ones. The corresponding alignment objective can be articulated as:
where denotes a truthful rejection, and an incorrect response. This suggests that refusal is preferable to providing an incorrect answer.
While this rejection strategy ensures reliability, it may unduly limit the model’s helpfulness. Therefore, we propose a more nuanced objective: the model should provide an answer whenever possible, categorizing its output as either high-confidence (“sure”) or low-confidence (“unsure”). The model must ensure high accuracy for “sure” responses while tolerating some errors in the “unsure” category, which can serve as references for users or inputs for further refinement. Our proposed alignment objective is formalized as:
Here, and denote correct and incorrect responses, while and represent “sure” and “unsure” confidence levels. This objective guarantees that correct “sure” predictions are prioritized over correct “unsure” predictions, which, while less confident, still provide utility. Incorrect “unsure” predictions are discouraged but considered more acceptable than incorrect “sure” predictions, as the latter significantly undermine overall reliability.
2.2 Reliability Evaluation
In this work, we focus on complex multi-slot problems and have modified several metrics for better evaluation of model reliability.
2.2.1 Weighted-F1
The F1-score, calculated using True Positives (TP), False Positives (FP), and False Negatives (FN), is typically the evaluation metric for multi-slot problems:
We modify the F1 score to align with the objectives of our EKBM framework. Since “unsure” predictions represent an state of uncertainty, we assign a weighted contribution (denoted by and , both ranging from 0 to 1) to reflect their partial utility as candidates for refinement. The modified precision and recall are defined as:
Here, STP and SFP denote True Positives and False Positives for “sure” predictions, while UTP and UFP are the corresponding values for the “unsure” category. For recall, we consider a prediction to be a UTP if it is partially correct (i.e., refinable). Unlike precision, we do not apply the weighting parameter to the denominator of the recall calculation, as recall measures the proportion of gold labels retrieved, which is independent of the introduction of . Notably, incorporating does not affect the range of precision and recall, both of which remain within [0,1]. The modified Weighted-F1 score is defined as:
2.2.2 Metric-Objective Consistency: Weighted-F1 as a Parametric Alignment Measure
The Weighted-F1 metric aligns with our four-tiered objective (Equation (3)) by using a parametric formulation that encodes hierarchical priorities, which ensures: (1) Strict enforcement for sure predictions: Unweighted terms prioritize high-confidence correctness (STP) and apply maximal penalties for errors (SFP). (2) Controlled utility for unsure predictions: The -scaled terms modulate low-confidence outputs, with attenuated UFP penalties tolerating unsure errors, and partial UTP rewards incentivizing refinable correct predictions.
This dual-parameter mechanism enables scenario-specific adaptation: Weighted-Precision regulates error tolerance via (higher tightens unsure-error constraints) and Weighted-Recall controls reference-value incentives via (higher promotes helpfulness through unsure-correct outputs). The composite metric thus operationalizes the reliability-helpfulness trade-off, permitting calibration from conservative (, ) to exploratory (, ) strategies.
We discuss a few corner cases: (1) = 0, = 0: Weighted-F1 disregards all “unsure” predictions. (2) = 0, = 1: Errors in the “unsure” category are ignored, rewarding effective recall, representing the theoretical upper limit after perfect refinement. (3) = 1, = 0: Lacks significant interpretation. (4) = 1, = 1: Eliminates confidence labeling, reverting to traditional methods.
2.2.3 Reliability Metrics
We use two specific Weighted-F1 scores: Quality-F1 as the primary metric for assessing model reliability, and Optimal-F1 as a supplementary reference.
Quality-F1: Defined as , this metric assigns half the weight to “unsure” predictions compared to “sure” ones. This encourages the model to provide high-confidence “sure” outputs while mitigating penalties for “unsure” ones. Quality-F1 directly measures the model’s self-awareness and reliability.
Optimal-F1 Defined as , it represents the theoretical upper limit achievable after perfect refinement of “unsure” predictions, reflecting the overall potential of the system.
3 Method
3.1 Explicit Knowledge Boundary Modeling
We formalize our proposed framework, Explicit Knowledge Boundary Modeling (EKBM), which aims to improve model reliability and self-awareness by explicitly modeling its knowledge boundaries and improve overall performance by incorporating a refinement model to improve the “unsure” predictions. As shown in Figure 2 (1), the framework divides the model’s decision-making process into two distinct stages, balancing immediate usability with comprehensive coverage, making the system both reliable and helpful. We provide a complete example of the two-stage execution flow and model outputs in Appendix B.
Fast Prediction with Confidence Labeling
In the first stage, the model performs the primary task and simultaneously assigns a confidence label (“sure” or “unsure”) to each prediction. “Sure” predictions are directly accepted as final outputs, ensuring high precision.
Slow Refinement for Unsure Prediction
In the second stage, “unsure” predictions undergo further refinement through strategies like user confirmation, automated multi-step reasoning, or other post-processing techniques, incuring extra cost but increases overall accuracy and reliability of the system.
3.2 Reliability Alignment
To align the model’s behavior with our alignment objective, we explore various training strategies to enhance its intrinsic self-awareness (Xu et al., 2025), including Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO). The data construction process is shown in Figure 2(2), with two detailed examples provided in Appendix §C.
SFT Data Construction
We assume that a model’s task-specific knowledge boundary can be approximated by repeated sampling on training data. Accordingly, our SFT data construction is straightforward: for each training sample, we perform multiple samplings. Predictions consistently deemed accurate are labeled “sure,” while those with errors or omissions are marked “unsure.” Crucially, we intentionally refrain from correcting erroneous outputs to avoid introducing supervisory signals that might inadvertently improve the model’s task performance. That is, when the model produces an incorrect prediction, we simply assign a confidence label of “unsure” without rectifying the error. The erroneous output remains in the final training data, reinforcing the model’s ability to recognize and categorize uncertain information rather than improving its accuracy directly.
The number of sampling rounds i is a hyperparameter. A larger i results in a more conservative model, as the proportion of “unsure” labels in the training corpus increases. In our study, we set i = 1 by default to accurately reflect the model’s realistic capability boundary.
DPO Preference Data Construction
DPO enables the model to learn directly from preference pairs, making the design of an effective preference strategy crucial. Given our alignment objective, it is intuitive to use Weighted-F1 as the preference score function. The values in Weighted-F1 significantly influence the model’s behavior and performance. Through experimentation (see §4.5), we determined that setting to 0.25 and to 0.75 best balances the model’s immediate helpfulness and its potential performance after refinement. Therefore, this setting will be used by default unless specified. To enhance the stability of DPO training, we ensure that all data is generated through multi-sampling from the model itself, aligning with the model’s output distribution.
3.3 Refinement Model
We train automated refinement models for each dataset to further optimize unsure predictions by multi-step reasoning, leveraging supervised fine-tuning (SFT) based on the LLAMA3 8B model. Our refinement operates at a fine-grained level, addressing each unsure prediction individually. We employ a reasoning paradigm similar to the Chain of Thought (Wei et al., 2022) methodology and use GPT-4o (Hurst et al., 2024) to generate detailed reasoning process for each training prediction. This approach is illustrated in Figure 2 (3), and detailed information can be found in Appendix §D.
4 Experiments
We conduct experiments focusing on three core research questions:
RQ1 Does our training pipeline outperform current methods in enhancing self-awareness?
RQ2 Does EKBM framework improve the overall performance of LLMs on complex tasks?
RQ3 Does our approach scale robustly across foundation models with varying capabilities?
4.1 Experiments Setup
4.1.1 Model and Baselines
We utilize the LLaMA3 8B model and the Qwen2.5 7B model as the backbone for our experiments. For reliability alignment, we conducted Reliability-SFT and integrated DPO to create two additional variants: Reliability-SFT+DPO Joint Training, which undergoes joint training with SFT and DPO, and Reliability-SFT+DPO Post Training, where DPO training follows initial SFT training, optimizing the model beyond the Reliability-SFT foundation. The key distinction lies in the different source distributions of preference data sampled for DPO. Training details could be found in Appendix §A.
To evaluate the reliability of our proposed models, we compare them against six baselines: three prompt-based and three uncertainty-based approaches (see Appendix §E). Notably, our focus is on the model’s self-awareness rather than task performance, so we do not include SOTA DST models. By default, we use only 1,000 samples for Reliability-SFT training and 2,000 for DPO, which is significantly less than traditional models.
Prompt-based
methods encompass three approaches: Direct, Verbose, and Self-Reflection (denoted as SR, Ji et al. (2023a)). The Direct approach generates predictions without confidence estimation, while Verbose produces both predictions and confidence labels simultaneously. SR assesses each prediction through self-evaluative reflection.
Uncertainty-based
methods include Token Probability (Prob, Manakul et al. (2023)), Self-Consistency (SC, Chen & Mueller (2024)), and P(True) (Kadavath et al., 2022). These methods use different techniques to classify outputs as “sure” or “unsure”. Prob calculates average token probabilities; SC measures the frequency of predictions across multiple samples; and P(True) determines the probability of the “True” token during model self-evaluation.
Method Type Method MultiWOZ-2.4 BiTOD SGD Opti. F1 Qual. F1 Opti. F1 Qual. F1 Opti. F1 Qual. F1 LLAMA3-8B (Dubey et al., 2024) Prompt Direct 72.48 77.98 73.38 73.38 73.37 64.83 65.39 65.39 26.46 28.71 25.12 25.12 Verbose 85.95 85.85 80.88 53.34 91.53 76.52 78.09 53.56 69.59 27.93 29.25 14.98 SR 86.43 81.34 81.34 61.43 92.39 76.32 78.84 48.60 65.32 33.61 33.37 21.30 Uncertainty Prob 80.05 80.66 77.74 69.03 83.50 72.44 73.40 60.30 36.93 32.28 27.74 22.86 SC 80.11 89.51 84.15 64.48 83.16 75.48 77.33 62.93 84.66 41.06 42.76 16.55 P(True) 72.82 79.00 73.50 70.81 72.44 66.26 65.97 63.06 24.78 29.93 26.26 26.24 Reliability (Ours) SFT 94.04 91.75 91.65 82.44 95.66 88.25 88.56 83.08 86.74 85.51 82.66 61.64 + DPO Joint 94.52 92.16 92.83 83.50 95.84 88.85 88.89 83.11 89.07 86.31 85.89 61.42 + DPO Post 93.59 93.10 91.75 83.55 95.17 88.00 88.35 85.78 82.26 83.47 79.70 64.14 Qwen2.5-7B (Yang et al., 2024) Prompt Direct 69.42 71.37 67.22 67.22 71.19 65.31 65.98 65.98 33.43 36.11 33.40 33.40 Verbose 83.02 71.14 69.05 58.88 78.62 77.22 75.23 64.78 65.78 31.80 32.13 21.74 SR 73.54 72.08 69.69 67.87 73.73 66.48 67.90 66.34 35.87 36.82 35.13 34.05 Uncertainty Prob 77.12 74.47 72.02 65.19 84.57 74.43 75.49 61.11 46.66 38.91 34.83 29.36 SC 83.33 90.20 84.67 57.00 87.65 74.83 77.42 51.43 71.47 61.08 53.38 25.22 P(True) 70.03 71.44 67.49 67.24 72.11 65.66 66.13 65.65 33.60 36.40 33.67 33.56 Reliability (Ours) SFT 94.16 91.93 91.98 82.03 96.08 89.13 89.33 83.01 90.69 83.86 83.09 64.63 + DPO Joint 95.29 92.88 93.03 80.86 95.99 89.86 89.51 82.00 91.23 85.78 83.95 64.48 + DPO Post 94.02 92.23 92.12 82.89 96.15 89.29 89.60 84.12 86.70 84.61 81.00 67.51
4.1.2 Dataset
We evaluate our methods on the Dialogue State Tracking (DST) task in task-oriented dialogue systems, which involves extracting slot-value pairs from multi-turn dialogues based on a predefined ontology (Xu et al., 2024a). We focus on the intricate multi-slot problem, where the model is likely to exhibit varying confidence levels for different parts of responses to the same query. Consequently, we do not adopt traditional QA or mathematical datasets. To ensure a robust evaluation, we utilize three datasets: MultiWOZ-2.4 (Ye et al., 2021), BiTOD (Lin et al., 2021) and SGD (Rastogi et al., 2020). For BiTOD, we adopt the English version, while for SGD, we randomly select 10,000 instances from the testset to mitigate excessive computational overhead. Common evaluation metrics for DST include Joint Goal Accuracy (JGA) and Slot-F1. For JGA, a sample is deemed correct only if all predictions are accurate without omissions, while Slot-F1 offers a slot-level assessment.
4.2 Reliability Evaluation
The experimental results on model reliability are summarized in Table 1. We evaluate our reliability-training pipeline against baseline algorithms, with a particular focus on the model’s self-awareness capabilities. For RQ1, our method significantly outperforms traditional prompt-based and uncertainty-based methods across various datasets, demonstrating a substantial improvement in self-awareness.
Our approach consistently achieves superior Quality-F1 scores, demonstrating improved ability to distinguish predictions as “sure” or “unsure” based on confidence. A high indicates reliable “sure” predictions, fostering user trust with minimal need for verification. Simultaneously, a high ensures the “unsure” category contains a broad range of potentially correct outputs, which forms a solid basis for refinement. Together, these metrics lead to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance. In essence, our system optimally balances precision and recall, providing both immediate reliability and a strong potential for post-refinement accuracy.
In our approach, DPO-based methods generally outperform SFT-only methods. Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, suggesting a more conservative stance that defers to subsequent refinement. In contrast, those undergoing Post Training yield superior Quality-F1 scores, indicating greater confidence and higher-quality predictions. This discrepancy arises because Joint Training samples from a pre-SFT distribution, which has poorer performance. This results in a dataset with a higher proportion of “unsure” predictions, leading to more cautious model behavior.
Notably, some baselines, such as Self-Consistency, show high , , and Optimal-F1 scores, suggesting initial superiority. However, this often indicates an overly conservative model that frequently labels predictions as “unsure.” This leads to a suboptimal balance between prediction types, as evidenced by a significantly lower Quality-F1 score. In an extreme case, a model could label all predictions as “unsure”, maximizing Optimal-F1 but shifting the entire burden to the refinement stage. This approach sacrifices immediate usability and fails to balance reliability and helpfulness.
4.3 Refinement for Unsure Prediction
As described in Section §3.3, we trained an automatic refinement model for each dataset. We refined the “unsure” predictions from the initial stage and merged them with the “sure” predictions to generate the final results. Table 2 shows the overall evaluation. For RQ2, our method achieves significantly better performance after refinement, with substantial improvements in Slot-F1 and JGA.
MultiWOZ-2.4 | BiTOD | SGD | |||||
| Slot-F1 | JGA | Slot-F1 | JGA | Slot-F1 | JGA | ||||
| LLAMA3-8B | |||||||||
| Prompt | Direct | 78.86 | 36.01 | 82.40 | 36.46 | 32.41 | 13.58 | ||
| Verbose | 76.80 | 36.69 | 76.93 | 27.99 | 23.11 | 9.00 | |||
| SR | 78.68 | 29.72 | 76.50 | 30.74 | 29.90 | 11.25 | |||
| Uncertainty | Prob | 75.89 | 24.19 | 73.88 | 24.86 | 28.48 | 9.63 | ||
| SC | 70.98 | 20.70 | 69.98 | 24.86 | 25.60 | 4.39 | |||
| P(True) | 73.88 | 19.18 | 67.07 | 19.03 | 26.29 | 7.34 | |||
SFT | 85.95 | 53.76 | 85.21 | 60.09 | 74.04 | 27.62 | ||
| + DPO Joint | 86.47 | 54.31 | 84.53 | 61.37 | 75.47 | 29.18 | |||
| + DPO Post | 86.55 | 56.33 | 85.87 | 64.95 | 73.96 | 28.13 | |||
| Qwen2.5-7B | |||||||||
| Prompt | Direct | 78.66 | 36.35 | 83.23 | 37.39 | 42.02 | 16.93 | ||
| Verbose | 67.59 | 26.45 | 74.37 | 27.38 | 27.79 | 9.19 | |||
| SR | 68.88 | 19.09 | 67.18 | 19.15 | 34.77 | 8.78 | |||
| Uncertainty | Prob | 74.18 | 23.56 | 73.97 | 25.64 | 35.91 | 10.81 | ||
| SC | 59.34 | 23.45 | 65.60 | 28.31 | 34.34 | 3.06 | |||
| P(True) | 67.52 | 17.05 | 66.25 | 17.99 | 33.66 | 8.07 | |||
SFT | 84.50 | 51.90 | 85.29 | 58.39 | 74.86 | 28.02 | ||
| + DPO Joint | 87.06 | 52.41 | 85.44 | 62.31 | 74.85 | 27.06 | |||
| + DPO Post | 87.57 | 53.19 | 85.86 | 63.61 | 74.95 | 28.51 | |||
Referring to §4.2, despite their higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement compared to Post Training models. This discrepancy stems from the limitations of the refinement model. Since the refinement model is not perfect, an excessive number of “unsure” predictions can lead to residual errors after refinement, incurring significant costs without proportional gains. This highlights the importance of a balanced “sure” and “unsure” classification. An ideal model should minimize unnecessary “unsure” predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
4.4 Scalability Analysis
To evaluate scalability, we experimented with foundation models of varying capabilities. We consider the original LLAMA3 8B as a low-performance baseline (denoted as Low), then fine-tuned two models with 1,000 (Medium) and 10,000 (High) samples to enhance their DST capabilities. We then continually conducted reliability training and evaluated the performance of the obtained reliable LLMs against baseline methods. As shown in Figure 3, our methods consistently outperform baselines after refinement, demonstrating strong scalability. This is particularly relevant for real-world scenarios where models are often domain-specifically tuned and possess high task capabilities. Detailed results in Appendix §F further support these findings.
As illustrated in Table 7 in Appendix §F, the advantage of DPO Post Training over Joint Training becomes more pronounced as foundation model performance improves. This is because our offline DPO algorithm’s limitations are exacerbated by higher-performing foundation models, negatively affecting the Joint Training data distribution and overall performance. Notably, in the SGD dataset, the benefits of DPO Joint Training diminish with increasing foundation model performance, even falling behind the SFT-only approach.
4.5 DPO Preference Strategies
We constructed our DPO preference data using Weighted-F1 as the ranking metric, as detailed in §2.2.1. Weighted-F1 incorporates two adjustable parameters: for Weighted Precision and for Weighted Recall. Using the Reliability-SFT trained LLAMA3 8B model, we generated DPO preference data with varying values and performed subsequent Post Training. Results on the MultiWOZ-2.4 dataset (Figure 4) show that the model achieves optimal performance at and , reaching near-best results at a relatively low cost.
Analysis indicates that increasing heightens sensitivity to errors in “unsure” predictions, reducing tolerance and the number of “unsure” outputs. Conversely, increasing amplifies rewards for successfully recalling “unsure” predictions, encouraging a more conservative stance. The optimal alpha settings depend on the specific scenario and refinement model performance; if the refinement model underperforms, the cost of excessive “unsure” predictions may not be justified. The flexible adjustment of both parameters allows for adaptation to diverse real-world scenarios.
4.6 Detailed Analysis
4.6.1 Prediction Types Analysis
To evaluate our model’s ability to make reliable confidence judgments, we further analyzed the distribution of its output categories and compared them to baselines. We examined five output categories: correct and incorrect predictions classified as “sure” or “unsure”, along with missing predictions. The results are presented in Figure 5.
Our findings highlight two key improvements: (1) Enhanced reliability: We observed a notable reduction in incorrect “sure” predictions, with most errors concentrated in the “unsure” category. This indicates a stronger self-assessment capability, allowing the model to better identify potential errors than baselines. Additionally, the “unsure” category consistently contains fewer correct predictions than incorrect ones, a distinction baseline models struggle to achieve. (2) Improved recall: Our approach significantly reduces missing predictions by capturing previously omitted relevant information within the “unsure” category. While some of these may still contain errors, this shift enhances overall helpfulness and provides a stronger foundation for refinement.
4.6.2 Cost Analysis
In this work, Cost refers to the additional overhead from the refinement process. We define the cost of refining an “unsure” prediction as a constant value of 1, equating Cost to the number of “unsure” predictions. For better comparison, we express Cost as the proportion of “unsure” predictions.
We compare the performance and Cost of various baselines in Figure 6. Results are shown for the MultiWOZ-2.4 dataset, with complete findings in Appendix §G. Our methods show significant performance improvements at a relatively low Cost. Additionally, as shown in Appendix Table 9, Cost decreases significantly as the foundation model’s performance improves, which is expected as a stronger model produces fewer uncertain predictions.
4.6.3 Refinement Model Comparison
We used various refinement models to enhance predictions from the LLAMA3 Reliability-SFT+DPO Post model. As shown in Table 7, we compared our trained models against GPT-4o and the DeepSeek series (DeepSeek-AI et al., 2025) (8B and 70B), all known for strong reasoning. Results show significant performance differences, with DeepSeek 8B underperforming while the 70B model nearly matches GPT-4o. Our task-specific fine-tuning with CoT substantially improves refinement accuracy.
4.7 Extension to Mathematical Problems
To validate the generalization of our method to different tasks, we conducted supplementary experiments on the widely used GSM8K mathematics dataset (Cobbe et al., 2021) using the Qwen2.5-7B-Instruct model with 1,000 training samples. We evaluated the model under two test settings: (a) with Chain-of-Thought (CoT) reasoning (Wei et al., 2023) and (b) by directly predicting the final answer. We tested both settings as LLMs generally perform differently on each. For the refinement model, we used the DeepSeek-R1-Distill-Llama-70B model, which has strong mathematical reasoning capabilities.
We adapted the F1-based evaluation metric from Section §2.2 to an accuracy-based definition:
As shown in Table 3, our method consistently outperformed both prompt-based and uncertainty-based baselines in improving the model’s self-assessment ability, achieving higher Quality-Accuracy. Furthermore, after applying refinement with the same refinement model, our approach yielded a higher final accuracy at a reasonable computational cost.
| Method Type | Method | Quality Accuracy | Final Accuracy | ||
| Prompt | Direct | 86.81 | 86.81 | ||
| Verbose | 90.8 | 90.98 | |||
| Uncertainty | Prob | 87.69 | 88.55 | ||
| SC | 89.41 | 91.48 | |||
SFT | 89.69 | 91.05 | ||
| + DPO Joint | 90.14 | 93.10 | |||
| + DPO Post | 92.73 | 95.30 |
| Method Type | Method | Quality Accuracy | Final Accuracy | ||
| Prompt | Direct | 19.94 | 19.94 | ||
| Verbose | 21.44 | 24.34 | |||
| Uncertainty | Prob | 24.49 | 76.50 | ||
| SC | 24.37 | 70.05 | |||
SFT | 27.31 | 77.94 | ||
| + DPO Joint | 26.16 | 78.24 | |||
| + DPO Post | 27.83 | 79.83 |
5 Related Work
The knowledge boundary of LLMs has emerged as a critical topic in recent research, highlighting their limitations in generating reliable outputs (Yin et al., 2024; Li et al., 2024b). Misalignment between model behavior and knowledge boundaries can result in factual hallucinations (Huang et al., 2023) and ambiguous responses (Liu et al., 2023).
Various methods have been proposed to assess these boundaries. Uncertainty-based approaches quantify prediction confidence through token probabilities (Manakul et al., 2023) and consistency (Chen & Mueller, 2024). Calibration strategies, including prompting (Tian et al., 2023) and fine-tuning (Tao et al., 2024) for improved confidence expression, align model confidence with prediction accuracy. Internal State Probing evaluates prediction factuality by analyzing model states, such as activations (Li et al., 2024a). While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes (Huang et al., 2024; Zhou et al., 2024; Pan et al., 2025; Shan et al., 2025).
To further align model behavior rather than merely estimate uncertainty, some studies guide models in self-reflection to minimize self-contradictions (Wei et al., 2022; Ji et al., 2023b), albeit with increased computational overhead. More recently, RL-based approaches leverage uncertainty estimation to train reward models, enhancing LLM truthfulness by rejecting queries beyond their capabilities (Xu et al., 2024b; Chen et al., 2024; Xue et al., 2024). While effective, these methods rely on the accuracy of uncertainty estimation algorithms and meanwhile may unintentionally reduce overall helpfulness.
6 Conclusion
In this work, we propose a two-stage framework that enhances the reliability of large language models by integrating fast and slow thinking paradigms. Through precise confidence assessment and high-accuracy refinement, we achieve a balance between reliability and usability. Extensive evaluations demonstrate that our method significantly improves model self-awareness and performance across complex tasks, establishing a new paradigm for enhancing the reliability of language models in error-sensitive applications.
Acknowledgments
This work is funded by the China NSFC Projects (62120106006, 92370206, and U23B2057) and Shanghai Municipal Science and Technology Projects (2021SHZDZX0102 and 25X010202846).
Appendix A Training Details
We modify the DeepSpeed-Chat (Yao et al., 2023) framework to accommodate our experimental setup. We adopt most of the default training parameters provided by DeepSpeed-Chat except that the learning rate for DPO training is set to 1e-7 and the maximum sequence length is set to 8096 on the SGD dataset and 2048 on the others. All models are trained for 1 epoch using NVIDIA A800 GPUs.
Appendix B Execution Example for Two Stages
We provide an example sample that includes the complete execution flow and model outputs for the two stages as shown in Table 4.
| Input | ||||||||
| Stage One: Reliable LLM Predictions | ||||||||
| Stage Two: Refinement for “Unsure” Predictions: | ||||||||
| Final Refined Predictions: | ||||||||
Appendix C Training Data Construction
We provide detailed examples to illustrate our training data construction process, including model inputs and the expected outputs the model is intended to learn. The data construction for SFT is outlined in Table 5, while that for DPO is detailed in Table 6.
| Input | |||||||||||
| Gold Labels | |||||||||||
| LLM Original Predictions | |||||||||||
| Constructed SFT Sample Output | |||||||||||
| Explanation | |||||||||||
| Input | |||||||||||
| Gold Labels | |||||||||||
| Multi-Sampling With the Foundation LLM | |||||||||||
| Constructed DPO Sample Output Pairs | |||||||||||
| Explanation | |||||||||||
Appendix D Refinement Model
We trained an automated refinement model for each dataset. The refinement model conducts in-depth step-by-step reasoning on unsure predictions based on task-specific criteria, operating within a defined action space. Our refinement models are built upon LLAMA3 8B utilizing SFT training with data constructed using OpenAI’s GPT-4o.
Action Space
The action space of the refinement model includes three actions: remove, reserve, and correct. A prediction should be reserved if entirely correct, should be corrected if partially correct and should be removed if entirely unnecessary. For any unsure predictions, a proper action from these three options is sufficient to achieve the required refinement.
Data Construction
We adopt a multi-step reasoning Chain-of-Thought (CoT) paradigm. The specific reasoning instructions and processes are detailed in Table 10. For data construction, we used OpenAI’s GPT-4o model. During the data generation process, we provided the model with the prediction to be refined and the corresponding golden label, then instructed it to generate a reasoning process based on predefined principles, as shown in Table 11. We combined the instruction, input, the prediction to be refined, the thinking steps, and the refinement result into a final training sample. For each dataset, we sampled 5,000 examples from the model’s unsure predictions as well as from randomly selected data in the training set, ensuring a balanced distribution of data across three actions.
Appendix E Baselines Details
Here we detail the specifics of the baselines.
Prompt-based
Direct: Utilizes prompts similar to those in Table 13 to guide the model in generating predictions directly, without incorporating confidence labels.
Verbose: Employs prompts akin to those in Figure 13 to direct the model, which outputs confidence labels (sure or unsure) alongside predictions based on its self-assessment.
Self-Reflection (SR): Builds on predictions from the Direct baseline, allowing the model to reflect on each prediction to assess its confidence level.
Uncertainty-based
Token Probability (Prob): Computes average token-level probabilities, classifying outputs as “sure” or “unsure” based on a threshold , with probabilities normalized and set to 0.8.
Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as “sure,” while others are marked as “unsure.” We conduct ten samples () and set to 50%.
P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding “True” than “False” as “sure,” otherwise as “unsure.”
Appendix F Scalability Analysis
The detailed results are presented in Table 7, demonstrating that our method maintains superiority across various capability baselines, achieving more reliable outcomes (higher , , and Quality-F1) and enhanced overall performance post-refinement (higher JGA).
Method Type Method MultiWOZ-2.4 BiTOD SGD Qual. F1 JGA Qual. F1 JGA Qual. F1 JGA 1000Samples-Trained LLAMA3 Based Prompt Direct 91.07 86.28 87.84 42.73 93.01 92.78 92.82 70.91 82.28 85.10 82.77 38.13 Uncertainty Prob 91.84 86.52 86.84 45.43 94.93 93.55 93.66 72.73 84.45 85.24 82.80 39.48 SC 92.68 92.37 84.31 43.18 95.70 95.12 92.05 70.37 86.61 92.30 78.71 28.62 Reliability (Ours) SFT 95.42 92.11 89.86 55.62 94.97 94.74 93.80 73.64 85.16 85.14 81.06 39.23 + DPO Joint 95.11 90.68 89.05 58.94 96.06 95.60 95.11 75.36 81.66 82.82 79.69 37.62 + DPO Post 96.33 92.15 89.91 59.12 96.22 95.54 95.43 77.70 86.67 85.44 83.27 41.48 10000Samples-Trained LLAMA3 Based Prompt Direct 94.51 93.30 93.63 63.19 95.35 95.14 94.34 80.76 83.24 83.75 83.11 43.96 Uncertainty Prob 95.77 94.06 92.95 67.39 95.95 95.53 95.23 81.26 84.37 84.61 83.08 43.52 SC 96.24 96.40 92.13 67.61 97.83 97.81 93.42 78.04 88.13 92.42 79.26 38.20 Reliability (Ours) SFT 96.98 95.86 94.40 70.85 96.44 96.20 96.03 80.54 85.30 83.92 83.57 45.11 + DPO Joint 96.98 95.61 94.71 70.84 97.15 96.65 95.96 82.35 82.31 82.36 80.75 40.48 + DPO Post 97.71 96.32 94.80 71.51 96.57 96.35 96.59 82.69 86.29 85.53 84.21 45.43
Appendix G Refinement Cost
The detailed overall results of cost are presented in Table 9 and 9. The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 8(a) and Figure 8(b), respectively, where cost refers to the proportion of unsure predictions.
Method Type Method MultiWOZ-2.4 BiTOD SGD LLAMA Qwen LLAMA Qwen LLAMA Qwen Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Verbose 87.28 30.62 80.69 38.05 90.30 74.07 SR 55.43 28.07 95.50 8.55 84.77 15.07 Uncertainty Prob 36.47 40.65 50.46 52.91 56.68 53.01 SC 69.26 58.91 58.05 62.59 96.80 92.71 P(True) 14.63 10.22 13.35 1.82 1.52 1.71 Reliability (Ours) SFT 32.67 30.98 20.03 21.78 51.24 55.96 + DPO Joint 31.96 31.98 21.19 23.63 57.72 45.29 + DPO Post 27.69 28.38 10.54 18.14 40.68 53.10
Method Type Method MultiWOZ-2.4 BiTOD SGD Medium High Medium High Medium High Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Uncertainty Prob 6.55 5.74 3.29 1.37 8.14 5.53 SC 28.93 12.17 6.25 2.12 41.68 41.42 Reliability (Ours) SFT 13.52 5.12 2.16 0.66 12.60 1.87 + DPO Joint 10.70 3.94 1.36 1.04 11.66 2.67 + DPO Post 13.32 8.59 3.44 1.46 19.01 4.19
| Refinement Examples | |||||||||||||||||||||||||
| GPT4o Prompt For Refinement Thinking Steps Generation | |||||||||||||||||||||||||||||||||
Appendix H LLM Instructions
The instructions used in this work are presented in this chapter, taking the MultiWOZ-2.4 dataset as an example. The instructions for the BiTOD dataset are similar.
H.1 Inference Instruction
The prompts used with the original untrained LLAMA3 model are shown in Tables 12 and 13. Among them, the Direct Inference Instruction requires the model to directly output the prediction results, while the Reliability Inference Instruction additionally requires the model to output a confidence label alongside the prediction.
| Direct Inference Instruction | |||||||||||||||||||||||
| Reliability Inference Instruction | |||||||||||||||||||||||||||||
H.2 Training Instruction
The prompts used with the trained LLAMA3 model are shown in Tables 14 and 15. The Direct Training Instruction is designed to train foundation models capable of directly outputting prediction results. In contrast, the Reliability Training Instruction is designed to train models capable of explicitly outputting confidence labels alongside predictions.
| Direct Training Instruction | |
| Reliability Training Instruction | |